can you help me out by helping me fine the center and radius of the circle
x^2+y^2-6x+4y-4=0

Answer by stanbon(48570)

find the center and radius of the circle
Complete the square separately onthe x-terms and the y-terms, as follows:
x^2+y^2-6x+4y-4=0
(x^2-6x+(6/2)^2) + (y^2+4y+(4/2)^2) = 4+(3^2)+(2^2)
(x+3)^2 + (y+2)2 = 17
Center at (-3,-2); radius = sqrt(17)
==============
Cheers,
Stan H.